# Robust image representations with counterfactual contrastive learning

### v1

@@ Introduction
SimCLR [9] is a widely adopted contrastive learning strategy, due to its effectiveness and simplicity in terms of training setup and number of hyperparameters to tune. Contrastive pairs are composed of two related views from the same image, obtained by applying a random augmentation pipeline to the original input. An image encoder then yields a high dimensional representation for each of the views, which is then projected onto a lower dimensional space using a two-layer perceptron to obtain representations $z$. The NT-Xent loss then pushes the representations of positive pairs closer together, while representations of negative pairs are pushed apart. For each positive pair $(i,j)$, the loss is:
$\mathcal{L}_{i,j}=-\log\frac{\exp(\text{sim}(z_{i},z_{j})/\tau)}{\sum_{k=1,k\neq i}^{2N}\exp(\text{sim}(z_{i},z_{k})/\tau)},$
where $\text{sim}(u,v)=\frac{u^{T}v}{||u||||v||}$.
@@
In this work, we propose a novel approach to contrastive positive pair creation, relying on domain counterfactuals instead of pre-defined random image transformations only. Applying this novel counterfactual pair creation pipeline to the classic SimCLR objective, we obtain ‘CF-SimCLR’, where we create positive view pairs by pairing each real image with one of its domain counterfactual. Concretely, we sample one target domain at random among all possible domains and generate the corresponding domain counterfactual to pair with the real image. If the original domain is sampled we simply keep the real image as the domain counterfactual (since there are no domain changes). To further increase view diversity, we then apply the original augmentation pipeline to this cross-domain positive pair. The rest of the SimCLR framework is kept as-is, as summarised in Fig. 1.
@@
Figure 2: Examples of counterfactual images generated with our model. Note that on PadChest, text is only imprinted on a subset of Imaging scans (not on Phillips): our model respects this by removing text when generating counterfactuals from Imaging to Phillips and vice-versa. Generated images have a resolution of 224x224 pixels for PadChest, 224x192 for EMBED.
Extension to other contrastive objectives: CF-DINO
The proposed counterfactual contrastive framework defines a novel way to create contrastive pairs by leveraging counterfactual image generation, independently of the particular choice of contrastive objective. To demonstrate that our counterfactual contrastive framework is indeed general and directly applicable to other contrastive learning objectives, we here extend our counterfactual contrastive analysis to models trained with the recently proposed and popular DINO-v2 objective [10]. As introduced in the related work section, DINO-v2 combines contrastive losses over different crops of images, model distillation and vision transformers to learn general image representations. For each image, we generate two ‘global views’ (i.e. larger crops of the images) as well as eight additional ‘local views’ (i.e. smaller crops). The model is then encouraged to produce similar representations for all global and local crops for both the student and teacher models. The final loss function is also complemented by a masked image modelling component. We invite the reader to refer to the original DINO-v2 paper for further details [10].
To incorporate our counterfactual contrastive strategy with DINO-v2, we follow similar steps as with CF-SimCLR. Specifically, in ‘CF-DINO’, one global crop is created from the real image while the other one is generated from its counterfactual. Similarly, we generate half of the local crops from the real image (N=4) and half from its counterfactual (N=4). During training, all views are encouraged to produce similar image representations, yielding the desired cross-domain invariance.
Experiments
Datasets
We evaluate the proposed method on two medical image modalities, mammography and chest radiography, using five public datasets covering a large variety of image acquisition hardware. The main chest radiography dataset used in this study is PadChest [30], a large dataset from Spain composed of scans acquired with two different scanners. In this dataset, scanner information is available for every image allowing us to easily train a domain counterfactual generation model. We use the same dataset for self-supervised pretraining. We evaluate the quality of the learned representations on pneumonia detection, first on in-distribution PadChest test data, then on two external datasets (covering acquisition domains unseen during pretraining): RSNA Pneumonia Detection [31, 32] and CheXpert [33]. For mammography, we primarily use the EMBED [34] dataset, containing over 300k scans, acquired in the US with 6 different devices. We keep one domain as hold-out domain (‘Senographe Essential’) and use the remaining five scanners for pretraining and counterfactual image generation. We highlight that 90% of the EMBED data was acquired with the ‘Selenia Dimensions’ scanner, the other scanners being heavily under-represented in this dataset, an ideal setup for investigating robustness to domain shifts. Finally, we investigate the quality of the learned encoders when transferring to the external VinDR-Mammo [35] dataset from Vietnam, covering two different acquisition domains. Table 1 details dataset splits and inclusion criteria.
Dataset Inclusion criteria Number of images
Train Validation Test
EMBED [34] 2D only(∗) 223,086 8,295 65,992
Senographe Essential [34] - 10,927 1,251 3,022
VinDR Mammo [35] - 11,191 2,813 5,996
PadChest [30] Adult PA only 64,989 7,203 17,993
CheXpert [33] PA only 13,811 2,449 10,838
RSNA Pneumonia [31] PA only 8,633 1,524 4,354
Table 1: Datasets splits and inclusion criteria. Splits are created at the patient level. (∗) excluding Senographe Essential, kept as separate hold-out domain.
Causal inference model
@@
To train the Deep Structural Causal Model, we need to specify the causal graph outlining the data-generating process. Most applications require the causal graph to closely describe true physiological and imaging processes affecting image appearance. However, in this study we only intervene on one variable, the ‘scanner’ indicator. As such we can simplify the causal graph to only contain this one variable, as factors of variations unaccounted for in the causal graphs will be captured by the exogenous noise. Importantly with this minimalist graph, we do not need to condition the generation model on any downstream task labels, essential to preserve the unsupervised nature of the pretraining step. In our examples, we include scanner as the only parent in the causal graph for mammography generation and include both biological sex and scanner for chest radiography counterfactual inference (sex is optional for domain counterfactual generation). We provide qualitative examples of generated domain counterfactuals in Fig. 2. Generated domain counterfactuals can deceive a domain classifier trained on real data 95% of the time for the PadChest model, and 80% of the time for the EMBED model (when generating counterfactuals uniformly across domains). For EMBED, we use weighted batch sampling during training to counter the imbalance in the scanner distribution.
Implementations details for pretraining and evaluation
@@
Pretraining. We use ResNet-50 [36] encoders (initialised with ImageNet weights) for all models pretrained with SimCLR. DINO-v2 use ViT-Base [15] encoders, initialised with the weights from ImageNet DINO-v2. We used the original DINO-v2 code and hyperparameters [15] for training and kept the encoder with the lowest validation loss. We kept all hyperparameters constant when comparing various contrastive pair generation strategies.
Evaluation. Models are evaluated with linear probing (i.e. classifiers trained on top of frozen encoders) as well as full model finetuning (unfrozen encoders) with varying amounts of annotated data. All models are finetuned with real data only, using a weighted cross-entropy loss and are evaluated on both ID and OOD datasets. Evaluation on external datasets is crucial to assess how counterfactual contrastive learning performs on unseen domains (outside of scanner distribution used for training the causal inference model). All our code is publicly available  [^1] https://github.com/biomedia-mira/counterfactual-contrastive.
Results

### v2

@@ 3.2 A simple and effective approach to counterfactual contrastive learning: CF-SimCLR
SimCLR (Chen et al., 2020) is a widely adopted contrastive learning strategy, due to its effectiveness and simplicity in terms of training setup and number of hyperparameters to tune. Contrastive pairs are composed of two related views from the same image, obtained by applying a random augmentation pipeline to the original input. An image encoder then yields a high dimensional representation for each of the views, which is then projected onto a lower dimensional space using a two-layer perceptron to obtain representations $z$. The NT-Xent loss then pushes the representations of positive pairs closer together, while representations of negative pairs are pushed apart. Concretely, for each positive pair $(i,j)$, the loss is given by:
$\mathcal{L}_{i,j}=-\log\frac{\exp(\text{sim}(z_{i},z_{j})/\tau)}{\sum_{k=1,k\neq i}^{2N}\exp(\text{sim}(z_{i},z_{k})/\tau)},$ $\text{sim}(u,v)=\frac{u^{T}v}{\|u\|\|v\|}.$ (1)
@@
In this work, we propose a novel approach to contrastive positive pair creation, relying on domain counterfactuals instead of pre-defined random image transformations only. Applying this novel counterfactual pair creation pipeline to the classic SimCLR objective, we obtain ‘CF-SimCLR’, where we create positive view pairs by pairing each real image with one of its domain counterfactuals. Concretely, we sample one target domain at random among all possible domains and generate the corresponding domain counterfactual to pair with the real image. If the original domain is sampled, we simply keep the real image as the domain counterfactual (since there are no domain changes). To further increase view diversity, we then apply the original augmentation pipeline to this cross-domain positive pair. The rest of the SimCLR framework is kept as-is, as summarised in Fig. 1.
The proposed method uses synthetically generated images for model training, in addition to the real images. Hence, the effective training size increases significantly compared to a model trained on real data only. As such, we need to disentangle the effect of the ‘smart pair creation’ proposed in CF-SimCLR and the effect of synthetically increasing the training set size. For this purpose, we here introduce another baseline, SimCLR+, where we add the same amount of synthetic examples in the training set as in CF-SimCLR. However, in this baseline, we do not implement our counterfactual pairing strategy: all images are considered independent samples, and standard contrastive learning is applied. As such, SimCLR+ and CF-SimCLR are trained using the exact same training set; only the positive pair creation mechanism differs. SimCLR+ is illustrated in green in Fig. 1.
3.3 Extension to other contrastive objectives: CF-DINO
The proposed counterfactual contrastive framework defines a novel way to create contrastive pairs by leveraging counterfactual image generation, independently of the particular choice of contrastive objective. To demonstrate that our counterfactual contrastive framework is indeed general and directly applicable to other contrastive learning objectives, we here extend our counterfactual contrastive analysis to models trained with the recently proposed and popular DINO-v2 objective (Oquab et al., 2023). As introduced in Section 2.1, DINO-v2 combines contrastive losses over different crops of images, model distillation and vision transformers to learn general image representations. For each image, we generate two ‘global views’ (i.e. larger crops of the images) as well as eight additional ‘local views’ (i.e. smaller crops). The model is then encouraged to produce similar representations for all global and local crops for both the student and teacher models. The final loss function is also complemented by a masked image modelling component. We invite the reader to refer to the original DINO-v2 paper for further details (Oquab et al., 2023).
To incorporate our counterfactual contrastive strategy with DINO-v2, we follow steps similar to those of CF-SimCLR. Specifically, in ‘CF-DINO’, one global crop is created from the real image while the other one is generated from its counterfactual. Similarly, we generate half of the local crops from the real image (N=4) and half from its counterfactual (N=4). During training, all views are encouraged to produce similar image representations, yielding the desired cross-domain invariance.
4 Experiments
4.1 Datasets
We evaluate the proposed method on two medical image modalities, mammography and chest radiography, using five public datasets covering a large variety of image acquisition hardware. The main chest radiography dataset used in this study is PadChest (Bustos et al., 2020), a large dataset from Spain composed of scans acquired with two different scanners. In this dataset, scanner information is available for every image, allowing us to train a domain counterfactual generation model easily. We use the same dataset for self-supervised pretraining. We evaluate the quality of the learned representations on pneumonia detection, first on in-distribution PadChest test data, then on two external datasets (covering acquisition domains unseen during pretraining): RSNA Pneumonia Detection (Stein et al., 2018; Shih et al., 2019) and CheXpert (Irvin et al., 2019). For mammography, we primarily use the EMBED (Jeong et al., 2023) dataset, containing over 300k scans, acquired in the US with 6 different devices. We keep one domain as a hold-out domain (‘Senographe Essential’) and use the remaining five scanners for pretraining and counterfactual image generation. We highlight that 90% of the EMBED data was acquired with the ‘Selenia Dimensions’ scanner, the other scanners being heavily under-represented in this dataset, an ideal setup for investigating robustness to domain shifts. Finally, we investigate the quality of the learned encoders when transferring to the external VinDR-Mammo (Nguyen et al., 2023) dataset from Vietnam, covering two different acquisition domains. Table 1 details dataset splits and inclusion criteria.
Dataset Inclusion criteria Number of images
Train Validation Test
EMBED 2D only(∗) 223,086 8,295 65,992
Senographe Essential - 10,927 1,251 3,022
VinDR Mammo - 11,191 2,813 5,996
PadChest Adult PA only 64,989 7,203 17,993
CheXpert PA only 13,811 2,449 10,838
RSNA Pneumonia PA only 8,633 1,524 4,354
Table 1: Datasets splits and inclusion criteria. Splits are created at the patient level. (∗) excluding Senographe Essential, kept as a separate hold-out domain.
4.2 Causal inference model and synthetic data generation
ScannerImage
(a) EMBED
ScannerSexImage
(b) PadChest
Figure 2: Causal graphs used to train the counterfactual image generation models used in this study.
@@ 4.2 Causal inference model and synthetic data generation
Figure 3: Examples of counterfactual images generated with our model. Note that on PadChest, text is only imprinted on a subset of Imaging scans (not on Phillips): our model respects this by removing text when generating counterfactuals from Imaging to Phillips and vice-versa. Generated images have a resolution of 224x224 pixels for PadChest and 224x192 for EMBED.
@@
To train the Deep Structural Causal Model, we need to specify the causal graph outlining the data-generating process. Most applications require the causal graph to closely describe true physiological and imaging processes affecting image appearance. However, in this study we only intervene on one variable, the ‘scanner’ indicator. As such, we can simplify the causal graph to only contain this one variable, as factors of variations unaccounted for in the causal graphs will be captured by the exogenous noise. Importantly, with this minimalist graph, we do not need to condition the generation model on any downstream task labels, which is essential to preserve the unsupervised nature of the pretraining step. In our examples, we include scanner as the only parent in the causal graph for mammography generation and include both biological sex and scanner for chest radiography counterfactual inference (sex is optional for domain counterfactual generation). We provide a visual representation of these causal graphs in  Fig. 2. For EMBED, we use weighted batch sampling during training to counter the imbalance in the scanner distribution. We provide qualitative examples of generated domain counterfactuals in Fig. 3. In terms of counterfactual effectiveness (Monteiro et al., 2022), generated scanner counterfactuals can deceive a scanner classifier trained on real data 94% of the time for the PadChest model, and 77% of the time for the EMBED model (when generating counterfactuals uniformly across domains). For sex counterfactuals on PadChest, we observed an effectiveness of 77% .
In CF-SimCLR, we construct positive pairs by combining a real image with a domain counterfactual, where the domain is randomly selected out of all possible domains available in the training set (see Section 3.2). Specifically, we first generate all possible domain counterfactuals using all scanners at hand in the training set. That is, for EMBED, for each image, we generate four possible counterfactuals (all available scanners except the real one); for PadChest, we generate one counterfactual for each image (the other scanner). This means that the combined training set of real + generated counterfactuals is balanced across scanners. We use this extended training set to train both CF-SimCLR and SimCLR+ (resp. CF-DINO and DINO+). The difference in scanner distribution between the original (real-only) training set and the counterfactual-augmented training set is depicted in Fig. 4.
(a) EMBED
(b) PadChest
Figure 4: Distribution of scanners in the original (real-only) training set and the counterfactual-augmented training set for EMBED (top) and PadChest (bottom).
4.3 Implementations details for self-supervised pretraining
@@
We use ResNet-50 (He et al., 2016) encoders (initialised with ImageNet weights) for all models pretrained with SimCLR. DINO-v2 use ViT-Base (Dosovitskiy et al., 2020) encoders, initialised with the weights from ImageNet DINO-v2. We used the original DINO-v2 code and hyperparameters (Dosovitskiy et al., 2020) for training and kept the encoder with the lowest validation loss. We kept all hyperparameters constant when comparing various contrastive pair generation strategies.
5 Results
In this section, we compare the quality and robustness of the learned representations for various pre-training paradigms. First, standard SimCLR. Secondly, SimCLR+, where we train a model using classic SimCLR on a training set enriched with domain counterfactuals. Finally, CF-SimCLR combines SimCLR with our proposed counterfactual contrastive pair generation framework. We then repeat the same analysis for models pre-trained with the DINO objective, comparing DINO, CF-DINO and DINO+. Note that in SimCLR+ (resp. DINO+), counterfactuals and real images are not paired during the contrastive learning step; they are all considered as independent training samples. As such, SimCLR+/DINO+ represent the common paradigm of simply enriching the training set with synthetic examples. In CF-SimCLR/CF-DINO, on the other hand, we systematically pair real images with their corresponding counterfactual for positive pair creation (Fig. 1).
We compare the effect of these three pretraining strategies on chest X-rays and mammograms. For chest X-rays, we evaluate the quality of the learned representations by assessing downstream performance on pneumonia detection. For mammography, we focus on the task of breast density prediction (important for risk modelling). Pre-training strategies are evaluated with linear probing (i.e. classifiers trained on top of frozen encoders) as well as full model finetuning (unfrozen encoders). Linear probes are best representative of the quality of learned representation during pre-training, as representations are unchanged during downstream training. However we also include the comparison with full model finetuning, as this is often the training paradigm of choice in practical scenarios. All models are finetuned with real data only, using a weighted cross-entropy loss. We evaluate the pre-trained encoders in two settings. First, we test the encoders on ID datasets, i.e. using the same data for pre-training, finetuning and testing. Secondly, encoders are evaluated on OOD datasets, i.e. where the model is finetuned/linear-probed and tested on data external to the pre-training data. Evaluation on external datasets is crucial to assess how counterfactual contrastive pretraining performs on unseen domains (outside of scanner distribution used for training the causal inference model). All encoders are pretrained on the full, unlabelled, PadChest and EMBED datasets. However, the main motivation for self-supervised pretraining is to increase robustness when only a limited amount of labelled data is available (Azizi et al., 2023). Hence, we evaluate the encoders for varying amounts of annotated data. Specifically, we finetune (or linear-probe) the encoder using a pre-defined amount of labelled samples, and then evaluate the resulting classifier on a fixed test set. We repeat this process several times, varying the amount of labelled samples to assess the effect of pre-training in function of number of labelled samples available for training the downstream classifier, e.g. for PadChest, the number of labelled training samples varies from 3,249 to 64,989. All our code is publicly available at https://github.com/biomedia-mira/counterfactual-contrastive.
For each task, we compare downstream performance across various pretraining strategies, for each scanner, for both SimCLR (Figs. 5 and 7) and DINO (Figs. 9 and 11) objectives. Moreover, to help visualise performance differences across encoders, we report the performance differences between the proposed encoders and the baseline in Figs. 6, 8, 12 and 10.
5.1 Does counterfactual contrastive learning improve performance and robustness under acquisition shift?
First, we focus on comparing counterfactual contrastive strategies (CF-SimCLR, CF-DINO) versus standard contrastive learning (SimCLR, DINO) to assess whether our cross-domain contrastive pair improve downstream performance across various domains (i.e. orange versus blue in Figs. 5 and 7 and orange bars in Figs. 6 and 8). We will compare CF-SimCLR with SimCLR+ in the next section.
Results on pneumonia detection, in Fig. 5, show that CF-SimCLR outperforms the SimCLR baseline (orange versus blue), across datasets, irrespective of the amount of labels, with improvements particularly striking for linear probing. Gains can be best observed on the performance difference plots in Fig. 6, where we observe that in all settings the performance difference compared to the SimCLR baseline is positive, i.e. CF-SimCLR performs better than SimCLR. For example, for linear probing, on PadChest, when using 3,249 training samples for training the Pneumonia classifier, the performance improves by 2.5% on the Imaging scanner from PadChest and 0.6% ROC-AUC on the Phillips scanner. On CheXpert, performance gains vary between 1% and 2% compared to the baseline with linear probing. Mammography results in Fig. 8 show that CF-SimCLR consistently outperforms the SimCLR baseline across most ID scanners for both linear probing and finetuning, particularly when the amount of labelled data is limited (<20k). On the external VinDR dataset, CF-SimCLR beats both baselines on both scanners in the low data regime. CF-SimCLR pretraining mainly benefits scanners under-represented in the training set (all except Selenia Dimensions for EMBED and PlanMed Nuance for VinDr), regardless of the amount of labelled data. For example, in linear probing, for EMBED, with 2,230 labelled samples, performance increases by 3% on Clearview CSM, 0.7% on Lorad Selenia, 1.3% on Pristina when comparing CF-SimCLR with SimCLR. Importantly, CF-SimCLR benefits the OOD scanners as well, in particular VinDR, where improvements range between 4%-6% ROCAUC compared to SimCLR when using 560-1,121 labelled samples. For higher levels of labels, the differences between encoders slowly vanish, particularly in finetuning settings. This is expected as starting representations matter less if many labels are available for training the classifier, during which learned representations may be substantially updated (in full model finetuning). For both modalities, the improvement on external datasets is particularly worth highlighting, given that the encoder was not exposed to these external domains during pretraining (nor during counterfactual generation).
Crucially, performance improvements and increased robustness to acquisition shift equally hold for encoders trained with the DINO-v2 objective, demonstrating the versatility of the proposed method. In Figs. 9 and 10, we can see that CF-DINO outperforms DINO for all EMBED scanners, across all levels of labels, for both linear probing and full model finetuning. Interestingly, on this dataset, we note that, for most scanners, the performance improvements between CF-DINO and DINO are even larger than between CF-SimCLR and SimCLR. Again, the gains mostly affect scanners under-represented during training with CF-DINO closing the performance gap between the majority scanner (Selenia Dimensions) and the other scanners (Fig. 9). For linear probes, in Fig. 10, we note that the improvement for low levels of labels (<20k) are substantial for some under-represented scanners on EMBED: on Senograph 2000D (resp. Clearview CSm) we see a 4% (resp. 2.8%) improvement with 2,223 labelled samples and 2% (resp. 1%) with 11,154 labels. On Lorad Selenia, improvements vary between 1% and 1.5% when training with up to 55k samples, whereas on Senographe Pristina, we observe 1% improvements with 2,223 labelled samples and 2% improvements for all other levels of labels. For VinDR, as for CF-SimCLR, performance improvements are mostly noticeable for the minority scanner PlanMed nuance, where performance gains are 5% for 560 labels, 6% for 1,121 labels and 2% for 2,803 labels, with gains vanishing for 11,212 labels (where all encoders perform equally). For finetuning, performance gains are mostly observed for low levels of labels and under-represented scanners (bottom two rows in Fig. 10). For chest X-rays, in Figs. 11 and 12, we can see that training with CF-DINO closes the performance gap between both PadChest scanners for both linear probing evaluation and finetuning. Indeed, without CF-DINO, there is a substantial performance drop from Phillips images to Imaging images, whereas this gap is much smaller with performance improving substantially with CF-DINO across all levels of labels for images from Imaging (Fig. 11). Similarly, in Fig. 12, we observed that CF-DINO outperforms DINO on the external RSNA Pneumonia dataset by a substantial margin 2% with 863 labelled samples, 3.5% with 2,158 labels and 1.1% with the full training set with linear probe, and improvements varying from 1 to 2% in finetuning as well. For CheXpert, CF-DINO outperforms DINO in model finetuning (and linear probing with 10% of labels). However, it slightly underperforms on the 25% and 100% for linear probing. Note that results on the CheXpert dataset are to be interpreted with some caution as labels are NLP sourced and, in general, of rather low quality (Irvin et al., 2019), which may explain the consequent drop of performance for all models between CheXpert and the expert-labelled RSNA Pneumonia dataset.
5.2 What is the benefit of counterfactual contrastive learning over simply extending the training set with the generated counterfactual data?
We have shown that the models trained with counterfactual contrastive learning outperform models trained with standard contrastive pairs. However, it is important to note that models trained with CF-SimCLR (resp. CF-DINO), are exposed to additional (synthetic) data during training. Hence, in this section, to isolate the effect of the ‘smart pair creation’ proposed in this work, we compare it to another baseline, SimCLR+ (resp. DINO+), where we use the same extended training set but where generated samples are considered as independent samples and are not paired with real images during training, details can be found in Section 3.2.
Differences between CF-SimCLR and SimCLR+ are best observed in Figs. 6 and 8 (resp. Figs. 12 and 10 for CF-DINO versus DINO+). Overall, CF-SimCLR outperformed SimCLR+ consistently across all experimental settings (Figs. 8 and 10), with similar results for CF-DINO. On mammography, CF-SimCLR/ outperformed SimCLR/DINO+ across all EMBED datasets. CF-SimCLR consistently also outperforms SimCLR+ on VinDr, while CF-DINO/DINO+ perform similarly on VinDR. On chest X-rays, in Fig. 6, CF-SimCLR consistently outperforms SimCLR+ across all settings. In Fig. 12, CF-DINO outperformed DINO+ by a large margin on PadChest and RSNA Pneumonia for linear probing, and both performed similarly for model finetuning. We especially noticed that the performance gains of SimCLR+ (resp. DINO+) over SimCLR were not very stable, improving for some domains while performing on par with standard SimCLR (resp. DINO) for others. In general, the counterfactual contrastive approaches offered much more consistent performance improvements. The starkest differences were observed on scanners under-represented during training (see Fig. 8).
These experimental results align with the theoretical benefits of counterfactual contrastive learning: by explicitly creating realistic cross-domain positive pairs during training, we directly encourage the network to create domain-agnostic image representations. The more domain-agnostic image representations are, the bigger the expected improvement in terms of robustness to acquisition shift. Such an increase in robustness leads to higher performance on devices under-represented during training, even more so in limited data settings. The increase in domain alignment is directly visible in the t-SNE plots of feature embeddings in Fig. 13, where embeddings of mammography models trained with SimCLR and SimCLR+ exhibit clear domain clustering, whereas the model trained with CF-SimCLR generate embeddings where domains are less separated. To further assess the domain separability of the features from those encoders, we train a scanner classifier on the features from each classifier. To reduce scanner-imbalance during training of the scanner classifier, we randomly select 1,000 images from the test set for each scanner (except for Senographe Pristina where we only had 101 test images), and extract corresponding features with each encoder. We then run 5-fold cross-validation, classifying scanners from the features of each pretrained model, using PCA (with 16 components) followed by logistic regression. Our results show that SimCLR and SimCLR+ achieve the same domain classification performance of 85% and 87% balanced accuracy, on the contrary CF-SimCLR only achieves 68% balanced accuracy, confirming that CF-SimCLR significantly reduces scanner separability in feature space.